# BLISS: A Lightweight Bilevel Influence Scoring Method for Data Selection in Language Model Pretraining

### v1

@@ Abstract
Effective data selection is essential for pretraining large language models (LLMs), enhancing efficiency and improving generalization to downstream tasks. However, existing approaches often require leveraging external pretrained models, making it difficult to disentangle the effects of data selection from those of the external pretrained models. In addition, they often overlook the long-term impact of selected data if the model is trained to convergence, primarily due to the prohibitive cost of full-scale LLM pretraining. In this paper, we introduce BLISS (BileveL Influence Scoring method for data Selection): a lightweight data selection method that operates entirely from scratch, without relying on any external pretrained oracle models, while explicitly accounting for the long-term impact of selected data. BLISS leverages a small proxy model as a surrogate for the LLM and employs a score model to estimate the long-term influence of training samples if the proxy model is trained to convergence. We formulate data selection as a bilevel optimization problem, where the upper-level objective optimizes the score model to assign importance weights to training samples, ensuring that minimizing the lower-level objective (i.e., training the proxy model over the weighted training loss until convergence) leads to best validation performance. Once optimized, the trained score model predicts influence scores for the dataset, enabling efficient selection of high-quality samples for LLM pretraining. We validate BLISS by pretraining 410M/1B/2.8B Pythia and LLaMA-0.5B models on selected subsets of the C4 dataset. Notably, under the 1B model setting, BLISS achieves $1.7\times$ speedup in reaching the same performance as the state-of-the-art method, demonstrating superior performance across multiple downstream tasks.
@@ 1 Introduction
The successful large-scale language model pretraining crucially relies on the careful choice of pretraining data (Brown et al., 2020; Raffel et al., 2020; Du et al., 2022; Elazar et al., 2023). Recent studies have shown that effective data selection (a.k.a., data curation) methods can enhance pretraining efficiency (Xie et al., 2023a) and improve generalization (Engstrom et al., 2024; Wettig et al., 2024). There are various types of data selection approaches for language model pretraining, including language filtering (Laurençon et al., 2022; Wenzek et al., 2019), data deduplication (Lee et al., 2021; Abbas et al., 2023), heuristic approaches (Rae et al., 2021; Penedo et al., 2023), data quality data filtering (Brown et al., 2020; Gao et al., 2020; Chowdhery et al., 2023; Xie et al., 2023b; Wettig et al., 2024), data mixing (Xie et al., 2023a; Albalak et al., 2023; Xia et al., 2023), and data influence function based methods (Park et al., 2023; Engstrom et al., 2024; Yu et al., 2024). Despite the rich literature of data selection methods in large language model (LLM) pretraining (e.g., a survey paper in Albalak et al. (2024)), it is still unclear what properties are needed for the training data curation to guarantee good performance: it remains an important real-world challenge (Li et al., 2024).
Existing approaches of data selection methods suffer from two major limitations. First, they often require leveraging pretrained models (Brown et al., 2020; Xie et al., 2023b; Wettig et al., 2024) for data-quality filtering, making it difficult to separate the effects of data selection from those of the external pretrained models. For example, the QuRating method (Wettig et al., 2024) assigns quality ratings to training samples based on responses from a pretrained LLM (e.g., GPT-3.5) before training a QuRater model. This reliance raises uncertainty about the role of the external LLM in the training process and whether its feedback is truly optimal. Moreover, the cost of invoking these external pretrained models is prohibitively expensive during data selection process for large-scale pretraining. Second, they typically do not consider the long-term impact of selected data if the model is trained to convergence. For example, the data influence function based approach (Yu et al., 2024) evaluates the impact of individual training samples based on a single training step with the current model, which does not capture the cumulative effects of data selection over the course of full model training.
@@
In this paper, we introduce a new data selection method, to address the two major limitations of existing approaches. Our method, namely BLISS (BileveL Influence Scoring method for data Selection), is a lightweight data selection method that operates entirely from scratch, without relying on any external pretrained models, while explicitly accounting for the long-term impact of selected data. The core innovation of our approach lies in the integration of two lightweight models within a novel bilevel optimization framework for data selection. Our method bypasses traditional data-quality filtering and explicitly considers the long-term impact of selected data throughout training. In particular, BLISS leverages a small proxy model as a surrogate for the LLM and employs a score model to estimate the long-term influence of training samples if the proxy model is trained to convergence. Our bilevel optimization problem has upper-level and lower-level objectives: the upper-level objective optimizes the score model to assign importance weights to training samples, ensuring that minimizing the lower-level objective (i.e., training the proxy model over the weighted training loss until convergence) leads to best validation performance. Once the bilevel optimization is solved, the trained score model predicts influence scores for the entire dataset, enabling the selection of high-score samples for LLM pretraining. The pipeline of our proposed procedure is illustrated in Figure 1. The main contributions of our paper are summarized as the following:
@@
  • We validate our method by pretraining 410M/1B Pythia and LLaMA-0.5B models on selected subsets of C4 dataset. Experimental results on 1B setting demonstrate a $1.7\times$ speedup in reaching the same performance as the state-of-the-art method such as MATES (Yu et al., 2024). Furthermore, we scale up our experiments to a 2.8B model pretraining used by the data selected in the 1B experiment, and we demonstrate that our method consistently outperforms MATES at every round of data selection, achieving $1.4\%$ performance improvement over MATES (Yu et al., 2024).
@@ 2 Related Work
Data Selection for Language Model Training. Early approaches to data selection primarily relied on rule-based methods as language filters for training data, employing utility functions tailored to specific datasets (Conneau & Lample, 2019; Raffel et al., 2020; Rae et al., 2021; Penedo et al., 2023). Another key category is data deduplication (Lee et al., 2021; Sorscher et al., 2022; Penedo et al., 2023; Abbas et al., 2023; Tirumala et al., 2023), which eliminates redundant samples to optimize training efficiency and enhance performance on downstream tasks. A class of methods exist for performing data-quality filtering, which can select data similar to high-quality corpus of data points (Brown et al., 2020; Du et al., 2022; Gao et al., 2020; Xie et al., 2023b), with small perplexity (Chowdhery et al., 2023; Wenzek et al., 2019). More recent methods leverage external pretrained LLMs to evaluate the pretraining data quality (Wettig et al., 2024; Maini et al., 2024). In addition, a similar variant of data selection is domain reweighting for data mixtures (Oren et al., 2019; Sagawa et al., 2019; Xie et al., 2023a; Fan et al., 2023; Albalak et al., 2023; Chen et al., 2024), which re-scale the contribution of each domain to enhance generalization. Another recently emerged line of research leverages the tool of influence functions (Hampel, 1974; Cook, 1977; Ling, 1984; Koh & Liang, 2017) to evaluate the impact of individual training samples on a fixed LLM (Park et al., 2023; Engstrom et al., 2024; Yu et al., 2024). In contrast to these works, our work explicitly considers the long-term impact of selected data if the model is not simply fixed but trained to convergence. In addition, our method can train the model from scratch and does not need any extra information from any external pretrained models, making it a scalable and effective solution.
@@
Bilevel Optimization and Data Selection. Bilevel optimization provides a powerful framework for modeling optimization problems with a nested structure (Bracken & McGill, 1973; Dempe, 2002). Recent research has focused on developing efficient bilevel optimization algorithms with strong theoretical guarantees (Ghadimi & Wang, 2018; Hong et al., 2023; Ji et al., 2021; Kwon et al., 2023; Dagréou et al., 2022; Chen et al., 2023; Grazzi et al., 2022; Hao et al., 2024; Gong et al., 2024). This approach has been widely applied in various machine learning tasks, including meta-learning (Finn et al., 2017), hyperparameter optimization (Franceschi et al., 2018), and natural language processing (Somayajula et al., 2023; Grangier et al., 2023). For the application of data selection, bilevel optimization has been utilized for continual learning (Borsos et al., 2020; Zhou et al., 2022; Hao et al., 2023) and data reweighting in LLM fine-tuning (Pan et al., 2024; Shen et al., 2024). Our work is most closely related to SEAL (Shen et al., 2024), which focuses on selecting high-quality and safe data to fine-tune a pretrained LLM, with the goal of aligning the model with safety and ethical guidelines. However, our approach differs from SEAL in two key aspects: (1) Problem setting. While SEAL operates in a fine-tuning context, our objective is to select data for pretraining an LLM from scratch, aiming to improve downstream performance without relying on any external pretrained models. (2) Model update mechanism. SEAL utilizes the LoRA technique (Hu et al., 2021) to update both the data selector and the LLM during fine-tuning. However, this approach is not directly applicable to our setting due to the following reasons. First, LoRA is only suitable for fine-tuning tasks but insufficient for full model pretraining. Second, their algorithm always updates the original large models directly, which is computationally expensive if all parameters are updated. In contrast, we propose a more efficient framework that introduces lightweight models (a score model and a proxy model) to guide data selection, while allowing full parameter updates within these smaller networks. To the best of our knowledge, our proposed bilevel influence scoring method is the first to leverage bilevel optimization techniques for data selection in LLM pretraining.
@@ 3 Preliminaries and Notations
Suppose that we have a large-scale training dataset $\mathcal{D}_{tr}=\{\xi_{i}\mid 0\leq i\leq N-1\}$ and a downstream task $\mathcal{D}_{ds}$. The goal is to select a subset of training set, namely $\mathcal{D}_{s}=\{\xi_{j}\mid 0\leq j\leq Q-1,Q\leq N\}$, to pretrain a large language model with a specific training budget (e.g., limited FLOPs), such that the model can achieve high performance on the downstream task $\mathcal{D}_{ds}$. Generally, the downstream data is inaccessible during pretraining. Instead, we can use a validation data $\mathcal{D}_{val}=\{\zeta_{i}\mid 0\leq i\leq M-1\}$ to estimate the model’s performance on $\mathcal{D}_{ds}$, because these two datasets often have similar data distributions or share common domain knowledge. A small subset of training data $\tilde{\mathcal{D}}_{tr}\subset\mathcal{D}_{tr}$ is uniformly sampled from $\mathcal{D}_{tr}$.
In bilevel optimization, $f(\cdot)$ and $g(\cdot)$ denote the upper-level and lower-level functions, respectively. Machine learning often requires solving stochastic optimization problems $f(\cdot)=\mathbb{E}_{\xi\sim\mathcal{D}_{f}}[F(\cdot;\xi)]$ and $g(\cdot)=\mathbb{E}_{\zeta\sim\mathcal{D}_{g}}[G(\cdot;\zeta)]$, where $\mathcal{D}_{f}$ and $\mathcal{D}_{g}$ are the underlying unknown data distribution for $f$ and $g$, respectively. Noisy observations of $f$ and $g$ can be collected by sampling from $\mathcal{D}_{f}$ and $\mathcal{D}_{g}$.
@@ 4.1 Bilevel Influence Scoring Framework
$\min_{\theta_{s}}\Phi(\theta_{s})$ $\coloneqq f(\theta_{p}^{*}(\theta_{s}))\coloneqq\mathbb{E}_{\zeta\sim\mathcal{D}_{\text{val}}}F\left(\theta_{p}^{*}(\theta_{s});\zeta\right),$ (1)
$\text{s.t. }\quad\theta_{p}^{*}(\theta_{s})$ $\in\operatorname*{arg\,min}_{\theta_{p}}g(\theta_{p},\theta_{s})\coloneqq\mathbb{E}_{\xi\sim\mathcal{D}_{tr}}G(\theta_{p},\theta_{s};\xi)$
where $\mathbb{E}_{\xi\sim\mathcal{D}_{tr}}G(\theta_{p},\theta_{s};\xi)=\sum_{i=0}^{N-1}P_{i}\mathcal{L}(\theta_{p};\xi_{i})+\gamma D_{KL}\left(\ell(\theta_{p};\xi_{i})\|\ell(\theta_{tr};\xi_{i})\right)+\lambda\|\theta_{p}\|^{2}$ and $P_{i}=\frac{e^{h(\theta_{s};\xi_{i})}}{\sum_{j=1}^{N}e^{h(\theta_{s};\xi_{j})}}$ represents the importance weight of sample $i$, and $h(\cdot):\mathbb{R}^{d_{x}}\rightarrow(0,1)$ is a function that maps a sample from $\mathbb{R}^{d_{x}}$ to an influence score in the range $(0,1)$. $\mathcal{L}(\cdot)$ and $F(\cdot)$ denote the loss functions for next token prediction, with a common choice being cross-entropy. The model’s output logits are represented by $\ell(\cdot)$. The KL divergence is defined as $D_{KL}(X\|Y)=\sum_{i}X_{i}\log(\frac{X_{i}}{Y_{i}})$. $\gamma$ and $\lambda$ are the regularization coefficients for the KL divergence and the weight decay terms, respectively.

### v3

@@ Abstract
Effective data selection is essential for pretraining large language models (LLMs), enhancing efficiency and improving generalization to downstream tasks. However, existing approaches often require leveraging external pretrained models, making it difficult to disentangle the effects of data selection from those of the external pretrained models. In addition, they often overlook the long-term impact of selected data if the model is trained for a long period of time, primarily due to the prohibitive cost of full-scale LLM pretraining. In this paper, we introduce BLISS (BileveL Influence Scoring method for data Selection): a lightweight data selection method that operates entirely from scratch, without relying on any external pretrained oracle models, while explicitly accounting for the long-term impact of selected data. BLISS leverages a small proxy model as a surrogate for the LLM and employs a score model to estimate the long-term influence of training samples if the proxy model is trained to convergence. We formulate data selection as a bilevel optimization problem, where the upper-level objective optimizes the score model to assign importance weights to training samples, ensuring that minimizing the lower-level objective (i.e., training the proxy model over the weighted training loss until convergence) leads to best validation performance. Once optimized, the trained score model predicts influence scores for the dataset, enabling efficient selection of high-quality samples for LLM pretraining. We validate BLISS by pretraining 410M/1B/2.8B Pythia and LLaMA-0.5B models on selected subsets of the C4 dataset. Notably, under the 1B model setting, BLISS achieves $1.7\times$ speedup in reaching the same performance as the state-of-the-art method, demonstrating superior performance across multiple downstream tasks.
@@ 1 Introduction
The successful large-scale language model pretraining crucially relies on the careful choice of pretraining data (Brown et al., 2020; Raffel et al., 2020; Du et al., 2022; Elazar et al., 2023). Recent studies have shown that effective data selection (a.k.a., data curation) methods can enhance pretraining efficiency (Xie et al., 2023a) and improve generalization (Engstrom et al., 2024; Wettig et al., 2024). There are various types of data selection approaches for language model pretraining, including language filtering (Laurençon et al., 2022; Wenzek et al., 2019), data deduplication (Lee et al., 2021; Abbas et al., 2023), heuristic approaches (Rae et al., 2021; Penedo et al., 2023), data-quality filtering (Brown et al., 2020; Gao et al., 2020; Chowdhery et al., 2023; Xie et al., 2023b; Wettig et al., 2024), data mixing (Xie et al., 2023a; Albalak et al., 2023; Xia et al., 2023), and data influence function based methods (Park et al., 2023; Engstrom et al., 2024; Yu et al., 2024). Despite the rich literature of data selection methods in large language model (LLM) pretraining (e.g., a survey paper in Albalak et al. (2024)), it is still unclear what properties are needed for the training data curation to guarantee good performance: it remains an important real-world challenge (Li et al., 2024).
Existing approaches of data selection methods suffer from two major limitations. First, they often require leveraging pretrained models (Brown et al., 2020; Xie et al., 2023b; Wettig et al., 2024) for data-quality filtering, making it difficult to separate the effects of data selection from those of the external pretrained models. For example, the QuRating method (Wettig et al., 2024) assigns quality ratings to training samples based on responses from a pretrained LLM (e.g., GPT-3.5) before training a QuRater model. This reliance raises uncertainty about the role of the external LLM in the training process and whether its feedback is truly optimal. Moreover, the cost of invoking these external pretrained models is prohibitively expensive during data selection process for large-scale pretraining. Second, they typically do not consider the long-term impact of selected data if the model is trained for a long time (i.e., multiple steps of gradient-based updates). For example, the data influence function based approach (Yu et al., 2024) evaluates the impact of individual training samples based on a single training step with the current model, which does not capture the cumulative effects of data selection over the course of full model training.
@@
In this paper, we introduce a new data selection method to address the two major limitations of existing approaches. Our method, namely BLISS (BileveL Influence Scoring method for data Selection), is a lightweight data selection method that operates entirely from scratch, without relying on any external pretrained models, while explicitly accounting for the long-term impact of selected data. The core innovation of our approach lies in the integration of two lightweight models within a novel bilevel optimization framework for data selection. Our method bypasses traditional data-quality filtering and explicitly considers the long-term impact of selected data throughout training. In particular, BLISS leverages a small proxy model as a surrogate for the LLM and employs a score model to estimate the long-term influence of training samples if the proxy model is trained to convergence. Our bilevel optimization problem has upper-level and lower-level objectives: the upper-level objective optimizes the score model to assign importance weights to training samples, ensuring that minimizing the lower-level objective (i.e., training the proxy model over the weighted training loss until convergence) leads to best validation performance. Once the bilevel optimization is solved, the trained score model predicts influence scores for the entire dataset, enabling the selection of high-score samples for LLM pretraining. The pipeline of our proposed procedure is illustrated in Figure 1. The main contributions of our paper are summarized as the following:
@@
  • We validate our method by pretraining 410M/1B Pythia and LLaMA-0.5B models on selected subsets of C4 dataset. Experimental results on 1B setting demonstrate a $1.7\times$ speedup in reaching the same performance as the state-of-the-art method such as MATES (Yu et al., 2024). Furthermore, we scale up our experiments to pretraining a 2.8B model using the data selected in the 1B experiment, and demonstrate that our method consistently outperforms MATES at every round of data selection, achieving $1.4\%$ performance improvement over MATES (Yu et al., 2024).
@@ 2 Related Work
Data Selection for Language Model Training. Early approaches to data selection primarily relied on rule-based methods as language filters for training data, employing utility functions tailored to specific datasets (Conneau and Lample, 2019; Raffel et al., 2020; Rae et al., 2021; Penedo et al., 2023). Another key category is data deduplication (Lee et al., 2021; Sorscher et al., 2022; Penedo et al., 2023; Abbas et al., 2023; Tirumala et al., 2023), which eliminates redundant samples to optimize training efficiency and enhance performance on downstream tasks. A class of methods exist for performing data-quality filtering, which can select data similar to high-quality corpus of data points (Brown et al., 2020; Du et al., 2022; Gao et al., 2020; Xie et al., 2023b; Li et al., 2024), with small perplexity (Chowdhery et al., 2023; Wenzek et al., 2019). More recent methods leverage external pretrained LLMs to evaluate the pretraining data quality (Wettig et al., 2024; Maini et al., 2024; Zhuang et al., 2025). In addition, a similar variant of data selection is domain reweighting for data mixtures (Oren et al., 2019; Sagawa et al., 2019; Xie et al., 2023a; Fan et al., 2023; Albalak et al., 2023; Chen et al., 2023b), which re-scale the contribution of each domain to enhance generalization. Another recently emerged line of research leverages the tool of influence functions (Hampel, 1974; Cook, 1977; Ling, 1984; Koh and Liang, 2017) to evaluate the impact of individual training samples on a fixed LLM (Park et al., 2023; Engstrom et al., 2024; Yu et al., 2024; Pan et al., 2025; Lin et al., 2024). QUAD (Zhang et al., 2024) proposes an efficient framework incorporating the attention layers to estimate the influence scores.
In contrast to these works, our work explicitly considers the long-term impact of selected data if the model is not simply fixed but trained for a long time. In addition, our method can train the model from scratch and does not need any extra information from any external pretrained models, making it a scalable and effective solution.
@@
Bilevel Optimization and Data Selection. Bilevel optimization provides a powerful framework for modeling optimization problems with a nested structure (Bracken and McGill, 1973; Dempe, 2002). Recent research has focused on developing efficient bilevel optimization algorithms with strong theoretical guarantees (Ghadimi and Wang, 2018; Hong et al., 2023; Ji et al., 2021; Kwon et al., 2023; Dagréou et al., 2022; Chen et al., 2023a; Grazzi et al., 2022; Hao et al., 2024; Gong et al., 2024). This approach has been widely applied in various machine learning tasks, including meta-learning (Finn et al., 2017), hyperparameter optimization (Franceschi et al., 2018), and natural language processing (Somayajula et al., 2023; Grangier et al., 2023). For the application of data selection, bilevel optimization has been utilized for continual learning (Borsos et al., 2020; Zhou et al., 2022; Hao et al., 2023) and data reweighting in LLM fine-tuning (Pan et al., 2024; Shen et al., 2024). Our work is most closely related to SEAL (Shen et al., 2024), which focuses on selecting high-quality and safe data to fine-tune a pretrained LLM, with the goal of aligning the model with safety and ethical guidelines. However, our approach differs from SEAL in two key aspects: (1) Problem setting. While SEAL operates in a fine-tuning context, our objective is to select data for pretraining an LLM from scratch, aiming to improve downstream performance without relying on any external pretrained models. (2) Model update mechanism. SEAL utilizes the LoRA technique (Hu et al., 2022) to update both the data selector and the LLM during fine-tuning. However, this approach is not directly applicable to our setting due to the following reasons. First, LoRA is only suitable for fine-tuning tasks but insufficient for full model pretraining. Second, their algorithm always updates the original large models directly, which is computationally expensive if all parameters are updated. In contrast, we propose a more efficient framework that introduces lightweight models (a score model and a proxy model) to guide data selection, while allowing full parameter updates within these smaller networks. To the best of our knowledge, our proposed bilevel influence scoring method is the first to leverage bilevel optimization techniques for data selection in LLM pretraining.
@@ 3 Preliminaries and Notations
Suppose that we have a large-scale training dataset $\mathcal{D}_{tr}=\{\xi_{i}\mid 0\leq i\leq N-1\}$ and a downstream task $\mathcal{D}_{ds}$. The goal is to select a subset of training set, namely $\mathcal{D}_{s}\subset\mathcal{D}_{tr}$ and $|\mathcal{D}_{s}|=Q\leq N$, to pretrain a large language model with a specific training budget (e.g., limited FLOPs), such that the model can achieve high performance on the downstream task $\mathcal{D}_{ds}$. Generally, the downstream data is inaccessible during pretraining. Instead, we can use a validation data $\mathcal{D}_{val}=\{\zeta_{i}\mid 0\leq i\leq M-1\}$ to estimate the model’s performance on $\mathcal{D}_{ds}$, because these two datasets often have similar data distributions or share common domain knowledge. A small subset of training data $\tilde{\mathcal{D}}_{tr}\subset\mathcal{D}_{tr}$ is uniformly sampled from $\mathcal{D}_{tr}$.
In bilevel optimization, $f(\cdot)$ and $g(\cdot)$ denote the upper-level (UL) and lower-level (LL) functions, respectively. Machine learning often requires solving stochastic optimization problems $f(\cdot)=\mathbb{E}_{\xi\sim\mathcal{D}_{f}}[F(\cdot;\xi)]$ and $g(\cdot)=\mathbb{E}_{\zeta\sim\mathcal{D}_{g}}[G(\cdot;\zeta)]$, where $\mathcal{D}_{f}$ and $\mathcal{D}_{g}$ are the underlying unknown data distribution for $f$ and $g$, respectively. $F(\cdot;\xi)$ denotes the upper-level stochastic objective function and $G(\cdot;\zeta)$ is the lower-level stochastic objective function. Noisy observations of $f$ and $g$ can be collected by sampling from $\mathcal{D}_{f}$ and $\mathcal{D}_{g}$.
@@ 4.1 Bilevel Influence Scoring Framework
$\min_{\theta_{s}}\Phi(\theta_{s})$ $\coloneqq f(\theta_{p}^{*}(\theta_{s}))\coloneqq\mathbb{E}_{\zeta\sim\mathcal{D}_{\text{val}}}F(\theta_{p}^{*}(\theta_{s});\zeta)\qquad\qquad\ \text{(UL)},$ (1)
$\text{s.t. }\theta_{p}^{*}(\theta_{s})$ $\in\operatorname*{arg\,min}_{\theta_{p}}g(\theta_{p},\theta_{s})\coloneqq\mathbb{E}_{\xi\sim\mathcal{D}_{tr}}G(\theta_{p},\theta_{s};\xi)\qquad\text{(LL)},$
where $\mathbb{E}_{\xi\sim\mathcal{D}_{tr}}G(\theta_{p},\theta_{s};\xi)=\sum_{i=0}^{N-1}P_{i}\mathcal{L}(\theta_{p};\xi_{i})+\gamma D_{KL}\left(\ell(\theta_{p};\xi_{i})\|\ell(\theta_{tr};\xi_{i})\right)+\lambda\|\theta_{p}\|^{2}$ and $P_{i}=\frac{e^{h(\theta_{s};\xi_{i})}}{\sum_{j=1}^{N}e^{h(\theta_{s};\xi_{j})}}$ represents the importance weight of sample $i$, and $h(\cdot):\mathbb{R}^{d_{x}}\rightarrow(0,1)$ is a function that maps a sample from $\mathbb{R}^{d_{x}}$ to an influence score in the range $(0,1)$. $\mathcal{L}(\cdot)$ and $F(\cdot)$ denote the loss functions for next token prediction, with a common choice being cross-entropy. The normalized model’s output logits are represented by $\ell(\cdot)$. The KL divergence is defined as $D_{KL}(X\|Y)=\sum_{i}X_{i}\log(\frac{X_{i}}{Y_{i}})$. $\gamma$ and $\lambda$ are the regularization coefficients for the KL divergence and the weight decay terms, respectively.